Government spending = R550 Exports = R330 Autonomous consumption = R280 Autonomous imports = R170 Investment expenditure = R120 Marginal propensity to consume = 0,75 Full employment level of income = R5700
What is the value of the multiplier?

[1] 2 [2] 2.50 [3] 4 [4] 3
1
Expert's answer
2020-04-15T10:01:56-0400

Level of the multiplier is (3)4

Because of


Multiplier= 1/mps=1/0.25 = 4


Multiplier=1/mps=1/0.25=4